1- ABSOLUTE VALUE OF (3-X) -5=2 (3-X)-5=2
2- X^3+2X^2-8X
3- 5X^2+8X-4
4- 6M^2-17M+12
5- 49X^2-25Y^2

Answer by Edwin McCravy(20056)

1.        |3 - X| - 5 = 2   
              |3 - X| = 2 + 5
              |3 - X| = 7

Break into two equations

      3 - X = 7  or  3 - X = -7
         -X = 4  or     -X = -10
         X = -4  or      X = 10

    Two solutions.  -4 and 10 
 


2.        X³ + 2X² - 8X
         X(X² + 2X - 8)
         X(X - 2)(X + 4)


3.       5X² + 8X - 4
         (5X - 2)(X + 2)


4.       6M² - 17M + 12
         (2M - 3)(3M - 4)

  
5.       49X² - 25Y²
         (7X - 5Y)(7X + 5Y)

Edwin